An aqueous solution containing 9.38 g of lead(II) nitrate is added to an aqueous solution containing 5.65 g of potassium chloride. Enter the balanced chemical equation for this reaction. Be sure to include all physical states. balanced chemical equation:

What is the limiting reactant?

The percent yield for the reaction is 90.4%. How many grams of precipitate is recovered? precipitate recovered:

How many grams of the excess reactant remain?
Answer #1

1)

Balanced chemical equation is:

Pb(NO3)2 (aq) + 2 KCl (aq) -> PbCl2(s) + 2 KNO3(aq)

2)

Molar mass of Pb(NO3)2,

MM = 1*MM(Pb) + 2*MM(N) + 6*MM(O)

= 1*207.2 + 2*14.01 + 6*16.0

= 331.22 g/mol

mass(Pb(NO3)2)= 9.38 g

use:

number of mol of Pb(NO3)2,

n = mass of Pb(NO3)2/molar mass of Pb(NO3)2

=(9.38 g)/(3.312*10^2 g/mol)

= 2.832*10^-2 mol

Molar mass of KCl,

MM = 1*MM(K) + 1*MM(Cl)

= 1*39.1 + 1*35.45

= 74.55 g/mol

mass(KCl)= 5.65 g

use:

number of mol of KCl,

n = mass of KCl/molar mass of KCl

=(5.65 g)/(74.55 g/mol)

= 7.579*10^-2 mol

Balanced chemical equation is:

Pb(NO3)2 + 2 KCl ---> PbCl2 + 2 KNO3

1 mol of Pb(NO3)2 reacts with 2 mol of KCl

for 2.832*10^-2 mol of Pb(NO3)2, 5.664*10^-2 mol of KCl is required

But we have 7.579*10^-2 mol of KCl

so, Pb(NO3)2 is limiting reagent

Answer: Pb(NO3)2

3)

Molar mass of PbCl2,

MM = 1*MM(Pb) + 2*MM(Cl)

= 1*207.2 + 2*35.45

= 278.1 g/mol

According to balanced equation

mol of PbCl2 formed = (1/1)* moles of Pb(NO3)2

= (1/1)*2.832*10^-2

= 2.832*10^-2 mol

use:

mass of PbCl2 = number of mol * molar mass

= 2.832*10^-2*2.781*10^2

= 7.876 g

% yield = actual mass*100/theoretical mass

90.4= actual mass*100/7.876

actual mass=7.12 g

Answer: 7.12 g

4)

According to balanced equation

mol of KCl reacted = (2/1)* moles of Pb(NO3)2

= (2/1)*2.832*10^-2

= 5.664*10^-2 mol

mol of KCl remaining = mol initially present - mol reacted

mol of KCl remaining = 7.579*10^-2 - 5.664*10^-2

mol of KCl remaining = 1.915*10^-2 mol

Molar mass of KCl,

MM = 1*MM(K) + 1*MM(Cl)

= 1*39.1 + 1*35.45

= 74.55 g/mol

use:

mass of KCl,

m = number of mol * molar mass

= 1.915*10^-2 mol * 74.55 g/mol

= 1.428 g

Answer: 1.43 g